IMACS progress report no. 59
B. Bigelow 20 March 2003
I) Project Status
Currently scheduled tasks: Status:
a) DSS drive assembly/testing (Hare, Kowal) Done
b) DSS drive test/integration (Birk) IP
c) Calibration sys. drafting (Cleary) Done
d) Calibration sys. fab (Kowal) IP
e) CCD array testing (Thompson/Burley) IP
f) Dewar/Detector final integration (Thompson/Burley) IP
g) Grism procurement (Sutin) IP
h) Instrument electronics assembly (Bagish) on sched.
i) Control software design (Birk) on sched.
j) Short Camera opto-assembly (Bigelow) Done
k) IFU design and build (U. Durham) IP
Project events in February:
- First light with short camera!
- Thermal testing delayed until final assembly complete
- Thermal testing enclosure test assembly completed
- IMACS pre-ship review delayed, currently unscheduled
- Magellan engineering time set aside for IMACS install - August '03
- Magellan FC/ACD optics delayed until September
Latest project images:
Final Assy: http://obs.carnegiescience.edu/instrumentation/imacs/devel/final/index.html
Disperser Server: http://obs.carnegiescience.edu/instrumentation/imacs/devel/dss/index.html
Filter Servers: http://obs.carnegiescience.edu/instrumentation/imacs/devel/fss/index.html
II) Optics (Sutin)
- Echellette prism due from Zygo 5/7/03
- Investigating calibration lamp types
- Started daily status report page (http://obs.carnegiescience.edu/~sutin/imacs/status.txt)
- Continued on-going flexure and optical testing
III) Electronics (Bagish)
- 3 power supply chassis assembled (guiders, array controllers)
- Need slide assembly for PS chassis to mount in racks
- L05 motor controller chassis finished and installed
- RS-422 communications chassis designed, panels in engraving
- Assembly of ABS box for PI controller in progress
- Cables ready for new IGUS chain
- Temp and pressure sensors ready for mounting
IV) Software (Birk/Clardy)
- Misc. motion control testing continues
- Problem identified with shutter motors - bad cable?
- Filter server testing finished except for discrete filter locations
- Improved optical modeling code - good to a pixel now.
- Created test files to continue mask cutter commissioning
- LDSS2 mask made with laser
- Design of IMACS mask making GUI due by 5/1/03
V) Detector Systems (Thomson/Burley)
- Working on CCD readout problems, bleeding
- Temp circuits added to pre-amp boards for additional temp sensors
- PCI and DSP PCB revisions in progress
- Working on temperature control code for CCDs
VI) Mechanics (Bigelow/Hare)
1) Finished short camera assembly and installed on instrument
2) Tested long camera focus motors with new PI controllers (w/ Birk)
3) Ordered all cooling system parts, started assembly of cooling system
4) Additional parts for new IGUS wrap and utility wrap designed, ordered
5) New getters designed, drafted, and fabricated for dewar
6) First grism cell fully assembled
7) Revised filter server insertion mechanisms completed and installed
1) Continue plumbing installation
2) Install second IGUS chain, re-route plumbing and cabling
3) Install first grism, both imaging mirrors on DSS
4) Continue final spacing measurements on collimator focus, cameras, dewar, BFD
5) Design shutter motor guards
6) Continue flexure testing
7) Re-mount S04 in new cell (from M&T)
- Continue assembly, testing, and rework of all guide cameras as necessary.
- Assembly and testing of Disperser Server System, including drive and
cable wrap sub-assemblies.
- Designed baffles for filter insertion stages
- Re-design center field guider stage to rotate barrel assemblies
VII) Fabrication (Kowal)
1) Finished machining components for all grating tilt mechanisms.
2) Made three getter assemblies.
3) New filter server arm & key made, plus additional modifications to FSS.
4) Various Dewar parts were made or modified.
5) Modified instrument rollers.
6) New parts for filter server made: spacers and top plates.
7) Cooling panels made.
8) Worked on disperser carousel, including: new limits switches, modifying
IGUS chains & correcting M & T manufacturing error on carousel.
Expectations for March:
1) Six grating cells complete.
2) Delrin filter server guides.
3) Parts for second imaging mirror support.
4) Calibration lamp reflector block.
5) Short camera barrel spacer.
6) Programming of cable tray parts.
IX) Other Business
Next meeting: 10:15 AM, 4/2/03